OpenStudy (anonymous):

I need to find the least common multiple of x^2-3x, x^2+3x, sorry can't write the problem right.. heh..

Parth (parthkohli):

Welcome on the site. Do you know how to factor each of them?

OpenStudy (nincompoop):

look at what you can use to be able to multiply \[x(x-3)\] x^2 + 3 doesn't have a common multiple unless you want to start using imaginary

OpenStudy (anonymous):

factored they are x(x-3) and x(x+3) not sure what to do from here

Parth (parthkohli):

Now write dem unique factors down in both your polynomials.

Parth (parthkohli):

I mean, take for example these two polynomials: \((x +1)(x - 3)\) and \((x + 3)(x + 1)\) The unique factors in both the polynomials are \(x + 1\), \(x - 3\) and \(x + 3\). Then multiply to get the lowest common multiple.

OpenStudy (anonymous):

so in x(x+3) and x(x-3) the unique ones are (x-3) and (x+3) and multiply them? so x^2-9?

OpenStudy (nincompoop):

x^2 + 3 doesn't have a common multiple if you are talking about elementary algebra.

Parth (parthkohli):

Brr, I read that as \(x^2 + 3x\) too. Silly me

OpenStudy (nincompoop):

imaginary is intermediate-level

OpenStudy (nincompoop):

oh lol!

OpenStudy (nincompoop):

in that case \[x(x+3)\]

OpenStudy (nincompoop):

I didn't know you can edit your question LOL

OpenStudy (anonymous):

lolsorry about that

OpenStudy (anonymous):

was that right? x^2-9?

Parth (parthkohli):

x, x+3 and x-3. Multiply

OpenStudy (nincompoop):

yes if you have NO middle term, then you might be looking at + and -

OpenStudy (anonymous):

x^2+3x(x-3)=x^3-3x+3x^2-9x= x^3+3x^2-12x?

OpenStudy (anonymous):

right?

OpenStudy (nincompoop):

what exactly are you doing?

OpenStudy (anonymous):

i don't even know anymore to be honest

OpenStudy (nincompoop):

\[x^3+3x^2-12 = x(x^2+3x-12)\]

OpenStudy (anonymous):

ah i see..

OpenStudy (anonymous):

thanks :)

OpenStudy (nincompoop):

try this link http://bit.ly/YFzI2k